40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 17 18:08:45 2017, Dhruva, Update, Optical Levers, Beam Profiling Setup 1.jpg2.jpg3.jpg4.jpg5.jpg
    Reply  Mon May 22 10:53:02 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  6x
       Reply  Tue May 23 10:27:24 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  6x
          Reply  Tue May 23 15:22:04 2017, rana, Update, Optical Levers, Beam Profiling Results  
             Reply  Tue May 30 18:31:54 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  plots.pdfspot_size_y.pdf
                Reply  Thu Jun 8 12:43:42 2017, Dhruva, Update, Optical Levers, Beam Profiling Results  profile.pdf
          Reply  Tue May 23 16:33:00 2017, Steve, Update, Optical Levers, Beam Profiling Results  
Message ID: 13008     Entry time: Tue May 23 16:33:00 2017     In reply to: 13006
Author: Steve 
Type: Update 
Category: Optical Levers 
Subject: Beam Profiling Results  

You may compare your results with this.

RXA: please no, that's not the right way

ELOG V3.1.3-